# Plasmon-plasmon interaction in nanoparticle assemblies: role of the dipole-quadrupole coupling

### v1

@@ Abstract
The synthesis of metallic nanoparticle assemblies is nowadays well-controlled, such that these systems offer the possibility of controlling light at a sub-wavelength scale, thanks, for instance, to surface plasmons. Determining the energy dispersion of plasmons likely to couple to light within these nanostructures is, therefore, a necessary preliminary task on the way to understanding both their photonic properties and their physical nature, namely the role of the quadrupole contribution. Starting with a general model that takes account of all energy modes, we show that its low-lying energy dispersion, gained numerically, can be compared to that of a minimal model that treats dipoles and quadrupoles on the same footing. The main advantage of the latter relies on the fact that its formulation is tractable, such that a semi-analytical Bogoliubov transformation allows one to access the experimentally relevant energy bands. Based on this semi-analytical derivation, we determine quantitatively the limit of validity of the dipole-only model, the presently proposed dipole and quadrupole model, compared to a full-plasmon-mode Hamiltonian. The results show that the dispersion relation, which includes dipoles and quadrupoles, is sufficient to capture the low-energy physics at play in most experimental situations. Besides, we show that at small lattice spacing, the contribution of quadrupoles is dominant around the Brillouin zone center.
@@ I Introduction
Plasmonic materials have proven to show a wide range of properties, with continuous research ranging from the fields of photonics [1], condensed matter [2], and spectroscopy. As a few striking examples, quantum entanglement was shown to survive a photon-plasmon-photon conversion [3], which justified a quantum description of the plasmons. A regime of very strong coupling with light was observed in plasmonic crystals, with the breakdown of the Purcell effect [4]. The possibility of investigating single-molecule vibrational states via plasmonic cavities was demonstrated [5, 6]. Among plasmonic materials, ordered plasmonic nanostructures are especially interesting as the control of their geometry enables the control of their optical properties [7, 8], and could lead to tailored properties of materials [9]. In the theoretical description of these superlattices, different models have been proposed. Some models use a classical description of the plasmons via point dipole [10, 11, 12] or point multipole models [13], and some use a quantum description of the plasmons via dipolar second quantization bosonic operators [14, 15, 16, 17] or multi-polar bosonic operators [18, 19, 20]. Regardless of the model chosen, it is always desirable to have a minimal model for the description of the physics at hand. That is why limiting the description to a coupled dipole model is often the chosen option. However, it is well established that this model stands for particles distant enough from one another. Indeed, a dipole-only model fails to describe the physics when the center-to-center distances ($d$) gets smaller than three times the radius ($d/R<3$) of the nanoparticles [21]. This is due to plasmon-plasmon induced hybridization [22], and it is observed even for small particles described in the quasistatic limit. This fact is particularly acute in recent interesting experimental realizations of colloidal crystals and other nanoparticle assemblies [4, 6], for which the nanoparticle center-to-center distance is well below the dipole model limit. Besides, considering coupling to fast electrons and their near fields requires higher order plasmons even for small particles, down to 4nm [23]. All these facts combine to suggest that a simple analytical expression of the low-energy plasmonic bands is necessary to help interpret future experimental data.
@@
In the present paper, we intend to take stock of this knowledge by building a minimal semi-analytic description, in line with earlier investigations, that necessarily goes, as seen above, beyond a simple dipole model. We thus investigate the importance of higher order plasmonic modes in dense assemblies. For the sake of clarity and simplicity, we choose a 1D chain of spherical nanoparticles as a model system, and we propose a minimal Hamiltonian leading to analytic expressions for the lowest energy plasmonic modes. Specifically, we discuss our derivation in light of previous similar approaches and show that taking account of the quadrupoles is sufficient, as it accurately describes the physics relevant to experiments.
@@ II Model
In the present work, we consider an assembly of metallic spherical nanoparticles embedded in a dielectric matrix as depicted, for instance, in Fig.1, for a 1D case. In this system, each nanoparticle hosts a discrete spectrum of localized surface plasmons $\left|\phi_{\ell,m}\right\rangle$ with corresponding energies $\hbar\omega_{\ell}$, where $\ell$ stands for the order of the mode and $m$ its azimuthal index. In second quantization, one can thus write a Hamiltonian $\sum_{\ell,m}\hbar\omega_{\ell}b_{\ell m}^{\dagger}b_{\ell m}$, where $b_{\ell m}^{\dagger}$ (respectively $b_{\ell m}$) creates (annihilates) a localized surface plasmon. For spherical particles, in the quasistatic limit, one has $\omega_{\ell}=\sqrt{\frac{\ell}{2\ell+1}}\omega_{p}$, where $\omega_{p}$ stands for the plasma frequency of the metal.
Within an assembly, the plasmon modes interact with one another, implying a hybridization, thereby forming bands. The plasmon-plasmon interaction terms can be evaluated either from overlap integrals $Q_{\ell m\ell^{\prime}m^{\prime}}$ [24, 25, 21] or from the Coulomb interaction in hydrodynamics models [22]. Taking this interaction into account, thanks to the overlap formulation, the plasmonic Hamiltonian reads
@@
The main advantage of this formulation relies on the fact that, in its general expression, the Eq. (2) allows one to describe any plasmonic nanoparticle assembly within the quasistatic limit, provided that the corresponding overlap term $Q_{\bm{k},\ell m\ell^{\prime}m^{\prime}}$ is correctly evaluated. To avoid unnecessary cumbersome expressions that would obscure the discussion, the present work limits its quantitative investigation to 1D systems.
The Hamilonian is block diagonal in $m$ since, for symmetry reasons, modes with different azimuthal indexes do not interact, as expressed in Eq. (5). In principle, as shown by Park and Stroud [21] a sharp description of the whole physics at play involves a large number of modes. However, if one focuses on the low-energy part of the spectrum, truncating at $\ell_{max}=2$ is sufficient, as shall be checked in Section III.1. In addition, from an experimental point of view this approximation appears reasonable: For instance, the plasmonic modes that will easily couple to light are $\ell=1$-dipolar modes and, as we shall see, those are strongly influenced by $\ell=2$-quadrupolar modes. Limiting oneself to $\ell_{max}=1$ in Eq. (2) leads to
@@
One can notice that this latter Hamiltonian is investigated by Allard and Weick in Ref.[14] and is restricted to dipolar modes. In the same spirit, but keeping in addition quadrupolar modes, we shall truncate the Hamiltonian in Eq. (2) at order $\ell_{max}=2$. This approximation corresponds to that of Ref.[19]. Yet, the derivation of the Hamiltonian by Barros et al., based on the development of canonically conjugate momenta for quadrupolar excitations, differs slightly from the present one and the expression is a bit more involved. Nevertheless, in the quasistatic limit, the two models come together and provide access to the same physics. Besides, if necessary, the present formulation allows one to investigate the role of higher-order multipoles, which is in contrast to Ref. [19], which is limited to quadrupoles.
@@
with $\ell,\ell^{\prime}=1,2$, enables to rewrite the Hamiltonian as
@@
The sum in the reciprocal space runs over the half-space $\mathcal{D}$. Indeed, one has to pay attention to the fact that in the definition of Eq. (8) the original bosonic operators carry either a positive or a negative $k$. The expressions for the commutator $[\mu_{\mathbf{k}\ell},H_{pl}]$ leads to solving $\det(M-\omega_{k}\mathds{1})=0$ with
@@ III.1 Superlattice spacing $d$
The dispersion of the different modes across the Brillouin zone is shown in Figure 3 for each case: $d/R=3$, $d/R=2.6$, and $d/R=2.4$. The converged $\ell_{max}=20$-model is plotted as gray and solid black lines, whilst the red dashed lines correspond to the dipole-model, and the blue crosses to the dipole+quadrupole model. The two low-lying highlighted bands are directly linked to dipolar modes. This is obvious if one looks at the $d/R=3$ case, where the three models produce nearly the same dispersion, and the deviation from a pure dipole model gets more pregnant as the inter-particle distance gets smaller. To understand the underlying mechanism, it is necessary to determine the nature of each band.
Figure 3: (a) Plasmonic band diagram of a chain of spherical metallic nanoparticles of radius $R=10$ nm with three different lattice spacing (a) $d=30nm$, (b) $d=26$nm and (c) $d=24$nm. Black lines are calculated with $\ell_{max}=20$, red dashed lines are the dipole model eigenvalues with $\ell_{max}=1$, and the blue crosses are the $\ell_{max}=2$ model taking into account quadrupoles. In (a), (b) and (c) the high energy bands calculated with $\ell_{max}=20$ are grayed out for readability. Insets show the corresponding chain densities.
The two highlighted bands in Fig. 3 correspond to the $m=0$ and $m=\pm 1$ indexes, the latter being two times degenerate owing to the symmetry of the chain. In the dipole model, the $m=0$ mode corresponds to dipoles aligned along the chain, while the $m=\pm 1$ relates to transverse modes with dipoles orthogonal to the chain axis. At the Brillouin zone center, the $m=0$-mode with all dipoles pointing in the same direction is the fundamental excitation, as it minimizes the dipolar tensor and can be viewed as a bonding mode. In contrast, at the edge $k=\pi/d$, a Néel-like order for the $m=1$-mode is lower in energy. This $\pi/d$ mode is referred to as an anti-bonding mode in the hybridization picture proposed by Nordlander et al. [22]. Therefore, a band-crossing occurs as one goes from $k=0$ to $\pi/d$, and this general observation does not depend on the superlattice spacing. However, the dipole model remains quantitatively valid only for $d/R\gtrsim 3$ but at smaller inter-particle spacings and finite $k$; this simple approach fails to provide an accurate description. The blue dashed line in Fig. 4 highlights this fact: it represents the relative difference $\Delta(k)=\left|\left(\omega(k)-\omega_{\ell=20}(k)\right)/\omega_{\ell=20}(k)\right|$ with respect to the converged $\ell_{max}=20$-model. In addition, this is further confirmed by the computation of the coefficient of determination $\mathsf{R}^{2}$
@@
Table 1: Values of $\mathsf{R}^{2}$ coefficient between the $\ell=20$ values and the truncated models values for the three lowest energy bands of different superlattices geometry corresponding to the plasmonic band diagrams of Figure 2.
In contrast, the dipole+quadrupole model, depicted by the blue crosses in Figs. 3 remains quantitatively valid down to $d/R=2.4$. As one can see in Fig. 4, truncating at $\ell_{max}=2$ leads to a relative energy difference that remains below $3\%$ across the whole $k$-space, even for dense assemblies and an $\mathsf{R}^{2}$, which is larger than $95\%$, comparable to that of the dipole model at $d/R=3$.
@@ III.2 Eigenstate analysis: role of the quadrupoles
For $d/R=3$, depicted at the top of Figs. 5 and 6, as expected, the dipolar component largely dominates the wave-function of Eq. (16) across the whole Brillouin zone. In the case of the $m=0$ band, and for both $d/R=2.6$ and $d/R=2.4$, the wave-function is purely dipolar at $k=0$, then it keeps a major dipolar contribution for small wave-vectors. At larger ones, a crossing occurs, and a pure quadrupolar state ends to take place at $k=\pi/d$. At high-symmetry points, quadrupoles and dipoles do not couple, which is in line with the results of Barros et al. [19], they only mix at finite wave-vectors and eventually cross. Even though the crossing wave-vector gets smaller as the lattice spacing is reduced, it remains in the order of $k_{{\rm cross}}\sim 0.6\pi/d$, which means that the corresponding modes will hardly couple to light
@@
Figure 5: (a) Components of the eigenfunctions associated with the band dispersion (12) for the $m=0$ band of a chain of spherical nanoparticles of radius $R=10$nm with three different lattice spacing (a) $d=30nm$, (b) $d=26$nm and (c) $d=24$nm. Red dashed lines are the dipole component $a_{\ell=1,m=0}(k)$, and the blue lines are the quadrupolar components $a_{\ell=2,m=0}(k)$
Figure 6: (a) Components of the eigenfunctions associated with the band dispersion (12) for the $m=\pm 1$ band of a chain of spherical nanoparticles of radius $R=10$nm with three different lattice spacing (a) $d=30nm$, (b) $d=26$nm and (c) $d=24$nm. Red dashed lines are the dipole component $a_{\ell=1,m=1}(k)$, and the blue lines are the quadrupolar components $a_{\ell=2,m=1}(k)$

### v2

# ]corresponding author

@@ Abstract
The synthesis of metallic nanoparticle assemblies is nowadays well-controlled, such that these systems offer the possibility of controlling light at a sub-wavelength scale, thanks, in parts to surface plasmons. Determining the energy dispersion of plasmons likely to couple to light in these nanostructures is, therefore, a necessary preliminary task on the way to understanding both their photonic properties and their physical nature, in particular the role of the quadrupole contribution. Starting with a general model that takes account of all energy modes, we show that its low-lying energy dispersion, gained numerically, can be compared to that of a minimal model that treats dipoles and quadrupoles on the same footing. The main advantage of the latter relies on the fact that its formulation is tractable, such that a semi-analytical Bogoliubov transformation allows one to access the experimentally relevant energy bands. Based on this semi-analytical derivation, we determine quantitatively the limit of validity of both the dipole-only model and the presently proposed dipole and quadrupole model, compared to a full-plasmon-mode Hamiltonian. The results show that the dispersion relation, which accounts for dipoles and quadrupoles, is sufficient to capture the low-energy physics in most experimental situations. Besides, we show that at small lattice spacing, the contribution of quadrupoles is dominant around the Brillouin zone center.
@@ I Introduction
Plasmonic materials have proven to show a wide range of properties, with continuous research ranging from the fields of photonics [1], condensed matter [2], and spectroscopy. As a few striking examples, quantum entanglement was shown to survive a photon-plasmon-photon conversion [3, 4], which justified a quantum description of the plasmons. A regime of very strong coupling with light was observed in plasmonic crystals, with the breakdown of the Purcell effect [5]. The possibility of investigating single-molecule vibrational states via plasmonic cavities was demonstrated [6, 7]. Among plasmonic materials, ordered plasmonic nanostructures are especially interesting as the control of their geometry enables the control of their optical properties [8, 9], and could lead to tailored properties of materials [10]. In the theoretical description of these superlattices, different models have been proposed. Some models use a classical description of the plasmons via point dipole [11, 12, 13] or point multipole models [14], and some use a quantum description of the plasmons via dipolar second quantization bosonic operators [15, 16, 17, 18] or multi-polar bosonic operators [19, 20, 21, 22]. The second quantized formalism becomes quite handy in describing the coupling with other fundamentally quantum systems, like molecules [23] or quantum dots [24] which makes it the formalism of choice here.
Regardless of the model chosen, it is always desirable to have a minimal model for the description of the physics at hand. That is why limiting the description to a coupled dipole model is often the chosen option. However, it is well established that this model stands for particles distant enough from one another. Indeed, a dipole-only model fails to describe the physics when the center-to-center distances ($d$) gets smaller than three times the radius ($d/R<3$) of the nanoparticles [25]. This is due to plasmon-plasmon induced hybridization [26], and it is observed even for small particles described in the quasistatic limit. This fact is particularly acute in recent interesting experimental realizations of colloidal crystals and other nanoparticle assemblies [5, 7], for which the nanoparticle center-to-center distance is well below the dipole model limit. Besides, considering coupling to fast electrons and their near fields requires higher order plasmons even for small particles, down to 4nm [27]. All these facts combine to suggest that a simple analytical expression of the low-energy plasmonic bands is necessary to help interpret future experimental data.
@@
In the present paper, we intend to take stock of this knowledge by building a minimal semi-analytic description, in line with earlier investigations. This description will necessarily go beyond a simple dipole model, as we have previously emphasized. We thus investigate the importance of higher order plasmonic modes in dense assemblies. For the sake of clarity and simplicity, we choose a 1D chain of spherical nanoparticles as a model system, and we propose a minimal Hamiltonian leading to analytic expressions for the lowest energy plasmonic modes. Specifically, we discuss our derivation in light of previous similar approaches and show that taking account of the quadrupoles is sufficient, as it accurately describes the physics relevant to experiments. The present minimal model implies tractable Hamiltonian matrices, and therefore combines two advantages: it allows one to derive a semi-analytical expression of the low-lying excitation dispersion, and the contribution of dipoles vs quadrupoles can be easily investigated. Besides, it sets the stage for a future coupling to photons, which is more straightforward in contrast to full multipole models.
@@ II Model
In the present work, we consider an assembly of metallic spherical nanoparticles embedded in a dielectric matrix as depicted, in Fig.1, for a 1D case. In this system, each nanoparticle hosts a discrete spectrum of localized surface plasmons $\left|\phi_{\ell,m}\right\rangle$ with corresponding energies $\hbar\omega_{\ell}$, where $\ell$ stands for the order of the mode and $m$ stands for its azimuthal index. In second quantization, one can thus write a Hamiltonian $\sum_{\ell,m}\hbar\omega_{\ell}b_{\ell m}^{\dagger}b_{\ell m}$, where $b_{\ell m}^{\dagger}$ (respectively $b_{\ell m}$) creates (annihilates) a localized surface plasmon. For spherical particles, in the quasistatic limit, one has $\omega_{\ell}=\sqrt{\frac{\ell}{2\ell+1}}\omega_{p}$, where $\omega_{p}$ stands for the plasma frequency of the metal. To give an order of magnitude, from experimental values, the bulk plasmon frequency of silver is $\hbar\omega_{p}^{{\rm Ag}}=9.04{\rm\ eV}$ and the one of gold is $\hbar\omega_{p}^{{\rm Au}}=8.89{\rm\ eV}$ [29]. These relatively high-energy values translate to the important role played in plasmonics by the nanoparticles’ size, as well as the surrounding matrix in assemblies [30].
Within an assembly, the plasmon modes interact with one another, implying a hybridization, thereby forming bands. The plasmon-plasmon interaction terms can be evaluated either from overlap integrals $Q_{\ell m\ell^{\prime}m^{\prime}}$ [28, 31, 25] or from the Coulomb interaction in hydrodynamics models [26]. Taking this interaction into account and using the overlap formulation the plasmonic Hamiltonian reads
@@
The main advantage of this formulation relies on the fact that, in its general expression, the Eq. (2) allows one to describe any plasmonic nanoparticle assembly within the quasistatic limit, provided that the corresponding overlap term $Q_{\bm{k},\ell m\ell^{\prime}m^{\prime}}$ is correctly evaluated. To avoid unnecessary cumbersome expressions that would obscure the discussion, the present work limits its quantitative investigation to 1D systems. Nevertheless, 2D and 3D situations can be addressed within the same framework. Yet, care has to be taken while computing the overlap of Eq. (5) from the real space expression in [25].
Owing to the symmetry of the chain, the present hamiltonian is block diagonal in $m$, and modes with different azimuthal indexes do not interact. In principle, as shown by Park and Stroud [25] a sharp description of the whole physics at play requires a large number of modes. However, if one focuses on the low-energy part of the spectrum, truncating at $\ell_{max}=2$ is sufficient, as shall be checked in Section III.1. In addition, from an experimental point of view this approximation appears reasonable: For instance, the plasmonic modes that will easily couple to light are $\ell=1$-dipolar modes and, as we shall see, those are strongly influenced by $\ell=2$-quadrupolar modes. Limiting oneself to $\ell_{max}=1$ in Eq. (2) leads to
@@
This latter Hamiltonian restricted to dipolar modes is investigated by Allard and Weick in Ref.[15]. In the same spirit, but keeping in addition quadrupolar modes, we shall truncate the Hamiltonian in Eq. (2) at order $\ell_{max}=2$. This approximation corresponds to that of Ref.[21].
Yet, the derivation of the Hamiltonian by Barros et al. is slightly different from the present one. It provides expressions of canonically conjugate momenta for quadrupolar excitations allowing for the coupling with a photonic bath, in the form of the quadrupolar tensor, and the expression is more involved. Nevertheless, in the quasistatic limit, the two models come together and provide access to the same physics. Besides, the present formulation allows one to investigate the role of higher-order multipoles, when Ref. [21], is limited to quadrupoles due once again to the difficulty of working out the coupling to free space electromagnetic field.
@@
where $u_{\mathbf{{k\ell}}},$ $v_{\mathbf{{k\ell}}},$ $w_{\mathbf{{k\ell^{\prime}}}}$ and $x_{\mathbf{{k\ell^{\prime}}}}$ represent a set of complex coefficients which are nothing but the elements of the unitary matrix involved in the Bogoliubov transformation. Their expression do not enter the computation of the energies conducted below from the commutators expressions. This expression with $\left(\ell,\ell^{\prime}\right)=\left(1,2\right)$, enables to rewrite the Hamiltonian as
@@
The sum in the reciprocal space runs over the half-space $\mathcal{D}$. Indeed, one has to pay attention to the fact that in the definition of Eq. (8) the original bosonic operators carry either a positive or a negative $k$. The expressions for the commutator $\mu_{\mathbf{k}\ell},H_{pl}]$ lead to solving $\det(M-\omega_{k}\mathds{1})=0$ with
@@ III.1 Superlattice spacing $d$
The dispersion of the different modes across the Brillouin zone is shown in Fig. 3 for each case: $d/R=3$, $d/R=2.6$, and $d/R=2.4$. For the sake of simplicity, one can take a formal example of $R=10{\rm\ nm}$, which is relevant to the quasistatic limit and provides interparticle distances of 24 nm in the $d/R=2.4$ case, in line with some typical colloidal crystals [32]. In fact, as the model is scale invariant, its limitation depends mainly on the quasistatic approximation validity. The converged $\ell_{max}=20$-model is plotted as gray and solid black lines, whilst the red dashed lines correspond to the dipole-model, and the blue crosses to the dipole+quadrupole model. The two low-lying highlighted bands are directly linked to dipolar modes. This is obvious if one looks at the $d/R=3$ case, where the three models produce nearly the same dispersion, and the deviation from a pure dipole model gets more pregnant as the inter-particle distance gets smaller. To understand the underlying mechanism, it is necessary to determine the nature of each band.
Figure 3: Plasmonic band diagram of a chain of spherical metallic nanoparticles of radius $R=10$ nm with three different lattice spacing (a) $d=30\ {\rm nm}$, (b) $d=26\ {\rm nm}$ and (c) $d=24\ {\rm nm}$. Black lines are calculated with $\ell_{max}=20$, red dashed lines are the dipole model eigenvalues with $\ell_{max}=1$, and the blue crosses are the $\ell_{max}=2$ model taking into account quadrupoles. The orange solid lines in (c) correspond to $\ell_{max}=3$ for comparison. In (a), (b) and (c) the high energy bands calculated with $\ell_{max}=20$ are grayed out for readability. Insets show the corresponding chain densities.
The two highlighted bands in Fig. 3 correspond to the $m=0$ and $m=\pm 1$ indexes, the latter being two times degenerate owing to the symmetry of the chain. In the dipole model, the $m=0$ mode corresponds to dipoles aligned along the chain, while the $m=\pm 1$ relates to transverse modes with dipoles orthogonal to the chain axis. At the Brillouin zone center, the $m=0$-mode with all dipoles pointing in the same direction is the fundamental excitation, as it minimizes the dipolar tensor and can be viewed as a bonding mode. In contrast, at the edge $k=\pi/d$, a Néel-like order for the $m=1$-mode is lower in energy. This $\pi/d$ mode is referred to as an anti-bonding mode in the hybridization picture proposed by Nordlander et al. [26]. Therefore, a band-crossing occurs as one goes from $k=0$ to $\pi/d$, and this general observation does not depend on the superlattice spacing. However, the dipole model remains quantitatively valid only for $d/R\gtrsim 3$ but at smaller inter-particle spacings and finite $k$ this simple approach fails to provide an accurate description. The blue dashed line in Fig. 4 highlights this fact: it represents the relative difference $\Delta(k)=\left|\left(\omega(k)-\omega_{\ell=20}(k)\right)/\omega_{\ell=20}(k)\right|$ with respect to the converged $\ell_{max}=20$-model. In addition, this is further confirmed by the computation of the coefficient of determination $\mathsf{R}^{2}$
@@
Table 1: Values of $\mathsf{R}^{2}$ coefficient between the $\ell=20$ values and the truncated models values for the three lowest energy bands of different superlattices geometry corresponding to the plasmonic band diagrams of Fig. 2.
In contrast, the dipole+quadrupole model, depicted by the blue crosses in Fig. 3 remains quantitatively valid down to $d/R=2.4$. As one can see in Fig. 4, truncating at $\ell_{max}=2$ leads to a relative energy difference that remains below $3\%$ across the whole $k$-space, even for dense assemblies and an $\mathsf{R}^{2}$, which is larger than $95\%$, comparable to that of the dipole model at $d/R=3$. Taking into account the higher-order $\ell=3$ octupolar modes corrects the model prediction only slightly as can be seen in Fig. 3, with $\mathsf{R}^{2}$ values of 0.99 at $d/R=2.4$ for both bands. In principle, the exact solution implies a Hamiltonian diagonalization with $\ell_{max}\to\infty$, it is therefore a matter of taste to truncate at an arbitrary mode once a convergence criterion is satisfactory. Of course, involving the octupoles implies a better quantitative match, but limiting ourselves to a dipole+quadrupole truncation has the advantage of providing a good physical description of the main low-lying excitations (within 5% quantitatively speaking). Besides, this approach allows one to obtain the tractable semi-analytical dispersion of Eq. (12), and provides a clear framework to understand the role of higher-order multipoles in plasmonic bands, as we shall see in the next Section.
As for the calculation of the plasmonic eigenvalues at a given $\bm{k}$, the dimensions of the Hamiltonian to be solved increase as $\sum_{\ell=1}^{\ell_{max}}(2\ell+1)=\ell_{max}^{2}+2\ell_{max}$. This gives Hamiltonian sizes of 3, 8, and 440 for the dipolar model, the quadrupolar model, and the model with $\ell_{max}=20$ respectively. From an experimental point of view, the main interest relies in the high symmetry $k=0$ point. To quantify the energy difference between the dipole model prediction, and the dipole+quadrupole model at this point, let us consider silver nanoparticles with $\hbar\omega_{p}=9.04{\rm\ eV}$. The dipole model’s predictions differ from those of the converged model $\ell_{max}=20$ by $\Delta E_{m=0}(k=0)=0.09{\rm\ eV}$ and $\Delta E_{m=1}(k=0)=0.29{\rm\ eV}$ while for the dipole+quadrupole model we have $\Delta E_{m=0}(k=0)=0.09{\rm\ eV}$ and $\Delta E_{m=1}(k=0)=0.05{\rm\ eV}$.
@@ III.2 Eigenstate analysis: role of the quadrupoles
For $d/R=3$, depicted at the top of Figs. 5 and 6, as expected, the dipolar component largely dominates the wave-function of Eq. (16) across the whole Brillouin zone. In the case of the $m=0$ band, and for both $d/R=2.6$ and $d/R=2.4$, the wave-function is purely dipolar at $k=0$, then it keeps a major dipolar contribution for small wave-vectors. At larger ones, a crossing occurs, and a pure quadrupolar state ends to take place at $k=\pi/d$. At high-symmetry points, quadrupoles and dipoles do not couple, which is in line with the results of Barros et al. [21], they only mix at finite wave-vectors and eventually cross. Even though the crossing wave-vector gets smaller as the lattice spacing is reduced, it remains in the order of $k_{{\rm cross}}\sim 0.6\pi/d$, which means that this would hardly couple to light.
@@
Figure 5: (a) Components of the eigenfunctions associated with the band dispersion (12) for the $m=0$ band of a chain of spherical nanoparticles of radius R=10R=10nm with three different lattice spacing (a) $d=30nm$, (b) d=26d=26nm and (c) d=24d=24nm. Red dashed lines are the dipole component $a_{\ell=1,m=0}(k)$, and the blue lines are the quadrupolar components $a_{\ell=2,m=0}(k)$
Figure 6: (a) Components of the eigenfunctions associated with the band dispersion (12) for the $m=\pm 1$ band of a chain of spherical nanoparticles of radius R=10R=10nm with three different lattice spacing (a) $d=30nm$, (b) d=26d=26nm and (c) d=24d=24nm. Red dashed lines are the dipole component $a_{\ell=1,m=1}(k)$, and the blue lines are the quadrupolar components $a_{\ell=2,m=1}(k)$